# Frontiers in Probabilistic Inference: Sampling Meets Learning

### v1

@@ 1 Introduction
Sampling from complex probability distributions is central to probabilistic inference and modern generative modeling. A recent line of work establishes dynamic measure transport as a unifying paradigm: construct a stochastic process $(X_{t})_{t\in[0,1]}$ whose marginals interpolate from a tractable source distribution $\mu_{0}$ to a target distribution $\mu_{1}$, estimate the conditional velocity $v_{t}(x)\coloneq\mathbb{E}[\dot{X}_{t}\mid X_{t}=x]$ and generate samples by evaluating the ODE flow maps $\phi_{t}$ defined by $\partial_{t}\phi_{t}(x)=v_{t}\!\left(\phi_{t}(x)\right)$ with initial condition $\phi_{0}(x)=x$. This perspective underlies methods such as stochastic interpolants ABVE (23); AVE (22), flow matching LCBH+ (23); TFM+ (24), and score-based probability flow ODEs SSDK+ (21), as well as rectified flows LGL (22); Liu (22); BRSR (25); HCD (25)—all of which have demonstrated strong empirical performance.
@@
Prior work (e.g., LGL (22); Liu (22); BRSR (25)) offers algorithmic frameworks that convert a given non-straight flow into a straight one, but a structural characterization of when straightness is intrinsic to the underlying process has remained open. This paper develops such a theory.
@@ 2.3 Linear characteristics
where $(X,Y)\sim\gamma$ and $\gamma$ is a coupling of the measures $\mu_{0}$ and $\mu_{1}$. Although seemingly restrictive, this process is widespread in computational statistics and generative modelling, e.g., optimal-transport displacement interpolation and modern flow-based generative modeling where linear sample-to-sample paths or linear noise–data blends define the training trajectory; see McC (97); LCBH+ (23); Liu (22); ABVE (23).
@@
$\nabla\cdot\left(\rho_{t}\,\Pi_{t}\right)=0\,.$
@@
Under the additional assumptions that $\mathbb{E}\,\|X_{t}\|^{2}<\infty$ and $\mathbb{E}\,\|\Pi_{t}(X_{t})\|^{2}<\infty$ for all $t\in[0,1]$, it holds that $\nabla\cdot\left(\rho_{t}\,\Pi_{t}\right)=0$ if and only if there exists a measurable map $T:D\to D$ satisfying $T_{\sharp}\mu_{0}=\mu_{1}$ such that $\left(\textup{id}\times T\right)_{\sharp}\mu_{0}=\gamma\,$, where $\textup{id}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the identity map.
@@ 3 Discussion
  • Full characterization of straightness. Derive necessary and sufficient conditions on the process $X$ under which the PDE (2) holds. Of particular interest is the case of $X_{t}=F(t,X_{0},X_{1})$ for a sufficiently regular $F:[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and a random vector $(X_{0},X_{1})\sim\mu_{0}\otimes\mu_{1}$. This is precisely the setting of stochastic interpolants AVE (22); ABVE (23) and readily leads to algorithmic insights.
@@ A.1 Notation
Let $(\Omega,\mathcal{F},\mathbb{P})$ be an abstract probability space. Random variables are measurable functions $X:\Omega\to\mathbb{R}^{d}$ [^1] In this paper, subsets of $\mathbb{R}^{d}$ are always equipped with the Borel $\sigma$-algebra. and the expectation $\mathbb{E}[X]$ is defined as the integral $\int_{\Omega}X\,d\mathbb{P}$. For random variables $X,Y:\Omega\to\mathbb{R}^{d}$ the conditional expectation $\mathbb{E}\left[X\mid Y\right]$ is defined at the conditional expectation $\mathbb{E}\left[X\mid\sigma(Y)\right]$ where $\sigma(Y)$ is the $\sigma$-algebra generated by $Y$ and for $y\in\mathbb{R}$ the conditional expectation $\mathbb{E}\left[X\mid Y=y\right]$ is the integral $\int_{\Omega}X\,d\mathbb{P}_{y}$ and $\mathbb{P}_{y}$ is the disintegration (Bog, 07, Section 10.6) of $\mathbb{P}$ on the level sets of $Y$. Finally, given another measurable space $(\Omega^{\prime},\mathcal{F}^{\prime})$ we denote the set of measures on $\Omega^{\prime}$ by $\mathcal{P}(\Omega^{\prime})$ and the subset of absolutely continuous measures by $\mathcal{P}_{\textup{a.c.}}(\Omega^{\prime})$.
@@ A.2 Proofs
if and only if there exists a measurable map $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ satisfying $T_{\sharp}\mu_{0}=\mu_{1}$ such that
@@
where $\textup{id}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the identity map.
@@
This concludes the proof of the “only if” part of the theorem. For the “if” part, notice that under the assumptions of the theorem we have $Y=T(X)$ almost surely and hence
$X_{t}=(1-t)\,X+t\,T(X)$
implying that $\textup{Var}(\dot{X}_{t}|X_{t}=x)=0$ for $\rho_{t}$-almost every $x\in\mathbb{R}^{d}$. This finishes the proof. ∎

### v2

@@ 1 Introduction
Sampling from complex probability distributions is central to probabilistic inference and modern generative modeling. A recent line of work establishes dynamic measure transport as a unifying paradigm: construct a stochastic process $(X_{t})_{t\in[0,1]}$ whose marginals interpolate from a tractable source distribution $\mu_{0}$ to a target distribution $\mu_{1}$, estimate the conditional velocity $v_{t}(x)\coloneq\mathbb{E}[\dot{X}_{t}\mid X_{t}=x]$ and generate samples by evaluating the ODE flow maps $\phi_{t}$ defined by $\partial_{t}\phi_{t}(x)=v_{t}\!\left(\phi_{t}(x)\right)$ with initial condition $\phi_{0}(x)=x$. This perspective underlies methods such as stochastic interpolants albergo2023stochastic; albergo2022building, flow matching lipman2022flow; tong2023cfm, and score-based probability flow ODEs song2021scorebased, as well as rectified flows liu2022flow; liu2022marginalpreserving; bansal2025wasserstein; hertrich2025relation—all of which have demonstrated strong empirical performance.
@@
Prior work (e.g., liu2022flow; liu2022marginalpreserving; bansal2025wasserstein) offers algorithmic frameworks that convert a given non-straight flow into a straight one, but a structural characterization of when straightness is intrinsic to the underlying process has remained open. This paper develops such a theory.
@@ 2.3 Linear characteristics
where $(X,Y)\sim\gamma$ and $\gamma$ is a coupling of the measures $\mu_{0}$ and $\mu_{1}$. Although seemingly restrictive, this process is widespread in computational statistics and generative modelling, e.g., optimal-transport displacement interpolation and modern flow-based generative modeling where linear sample-to-sample paths or linear noise–data blends define the training trajectory; see mccann1997convexity; lipman2022flow; liu2022marginalpreserving; albergo2023stochastic.
@@
$\nabla\cdot\left(\rho_{t}\,\Pi_{t}\right)=0\,.$ (3)
@@
Under the additional assumptions that $\mathbb{E}\,\|X_{t}\|^{2}<\infty$ and $\mathbb{E}\,\|\Pi_{t}(X_{t})\|^{2}<\infty$ for all $t\in[0,1]$ the equation $\nabla\cdot\left(\rho_{t}\,\Pi_{t}\right)=0$ implies that there exists a measurable map $T:D\to D$ satisfying $T_{\sharp}\mu_{0}=\mu_{1}$ such that $\left(\textup{id}\times T\right)_{\sharp}\mu_{0}=\gamma\,$, where $\textup{id}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the identity map. Moreover, if there exists such a continuously differentiable map $T$ with Jacobian $\nabla T$ having no zero or negative singular values, then equation (3) holds.
@@ 3 Discussion
  • Full characterization of straightness. Derive necessary and sufficient conditions on the process $X$ under which the PDE (2) holds. Of particular interest is the case of $X_{t}=F(t,X_{0},X_{1})$ for a sufficiently regular $F:[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and a random vector $(X_{0},X_{1})\sim\mu_{0}\otimes\mu_{1}$. This is precisely the setting of stochastic interpolants albergo2022building; albergo2023stochastic and readily leads to algorithmic insights.
@@ A.1 Notation
Let $(\Omega,\mathcal{F},\mathbb{P})$ be an abstract probability space. Random variables are measurable functions $X:\Omega\to\mathbb{R}^{d}$ [^1] In this paper, subsets of $\mathbb{R}^{d}$ are always equipped with the Borel $\sigma$-algebra. and the expectation $\mathbb{E}[X]$ is defined as the integral $\int_{\Omega}X\,d\mathbb{P}$. For random variables $X,Y:\Omega\to\mathbb{R}^{d}$ the conditional expectation $\mathbb{E}\left[X\mid Y\right]$ is defined at the conditional expectation $\mathbb{E}\left[X\mid\sigma(Y)\right]$ where $\sigma(Y)$ is the $\sigma$-algebra generated by $Y$ and for $y\in\mathbb{R}$ the conditional expectation $\mathbb{E}\left[X\mid Y=y\right]$ is the integral $\int_{\Omega}X\,d\mathbb{P}_{y}$ and $\mathbb{P}_{y}$ is the disintegration (bogachev2007measure, Section 10.6) of $\mathbb{P}$ on the level sets of $Y$. Finally, given another measurable space $(\Omega^{\prime},\mathcal{F}^{\prime})$ we denote the set of measures on $\Omega^{\prime}$ by $\mathcal{P}(\Omega^{\prime})$ and the subset of absolutely continuous measures by $\mathcal{P}_{\textup{a.c.}}(\Omega^{\prime})$.
@@ A.2 Proofs
Corollary 11.
For all $(t,x)\in[0,1]\times D$ we have:
$D_{t}v_{t}=0\,\iff\,\nabla\cdot\left(\rho_{t}\,\Pi_{t}\right)=\rho_{t}\,a_{t}\,.$
@@
implies that there exists a measurable map $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ satisfying $T_{\sharp}\mu_{0}=\mu_{1}$ such that
@@
where $\textup{id}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the identity map. Moreover, if there exists such a continuously differentiable map $T$ with Jacobian $\nabla T$ having no zero or negative singular values, then it holds that
$\nabla\cdot\left(\rho_{t}\,\Pi_{t}\right)=0\,.$
@@
This concludes the first part of the proof. For the second part, we follow the proof of (marzouk2025approximation, Theorem 3.4). Using the assumptions on the map $T$ we have that the map
$G:(x,t)\mapsto\big((1-t)\,x+t\,T(x),t\big)\,$,
is a bijection onto its image. Therefore, there is an inverse map $S:\textup{im}\,G\to\mathbb{R}^{d}$ such that
$S\big((1-t)\,x+t\,T(x),t\big)=x\,$,
which allows us to conclude that for each $t\in[0,1]$ we have
$S(X_{t},t)=X\text{ a.s. }$
Finally, recall that
$\dot{X}_{t}=T(X)-X\text{ a.s. }$
and thus
$\dot{X}_{t}=T\big(S(X_{t},t)\big)-S(X_{t},t)\text{ a.s. }$
showing that at each $t\in[0,1]$ the random variable $\dot{X}_{t}$ is a measurable function of $X_{t}$. This implies that $\textup{Var}(\dot{X}_{t}|X_{t})\equiv 0$ almost surely, completing the proof. ∎